OpenStudy (anonymous):

I GIVE MEDALS: Solve for m: 3m - 5r = 2n + 7 options A. (2n + 7 + 5r)/3 B. (2n + 7 - 5r)/3 C. m = (5r)/3 (2n + 7) D. m = (3)/(5r) (2n + 7)

OpenStudy (anonymous):

c

OpenStudy (anonymous):

wrong answer

OpenStudy (anonymous):

really???

OpenStudy (anonymous):

which variable are you solving for?

OpenStudy (aum):

3m - 5r = 2n + 7 To solve for m, isolate m. Add 5r 3m = 2n + 5r + 7 divide by 3 m = (2n + 7 + 5r) / 3

OpenStudy (anonymous):

ofcourse its not c its 'a' \[3m-5r=2n+7\] \[3m=2n+7+5r\] \[m=(2n+7+5r)3\]